OpenStudy (anonymous):

sin(x/7)sin(x/7) = 1 [ 1-cos(2x/7) ] true or false

OpenStudy (anonymous):

false

OpenStudy (anonymous):

false

OpenStudy (anonymous):

its true ...good job guys

OpenStudy (anonymous):

ty for another wrong answer openstudy ....thats what ..4 in a row now?

OpenStudy (anonymous):

tbh ..don't really expect medals on this one ..bye

OpenStudy (anonymous):

no its false

OpenStudy (anonymous):

No...its true. If it was false i would have gotten the question right now wouldn't i have.

OpenStudy (anonymous):

That would be true if sin(x/7)sin(x/7) = 1/2 [ 1 - cos(2x/7)] First know what you are typing...

OpenStudy (anonymous):

I typed it bc it said 1 ._.

OpenStudy (anonymous):

sinx/7 sin x/7=sin^2 x/7= 1-c0s^2x/7

OpenStudy (anonymous):

sin^2(theta) = [1 - cos(2theta)] / 2